# Scoring, Reasoning, and Selecting the Best! Ensembling Large Language Models via a Peer-Review Process

### v1

@@ 2.2.1 Graphical Model
where the distribution is parametrized by $\bm{\alpha}$. Next, similar to the concept of confusion matrix commonly used in machine learning (Bishop & Nasrabadi, 2006; Goodfellow, 2016), we introduce an annotator-specific transition matrix $\bm{\Pi}^{(j^{\prime})}$ to model the probability that an LLM confuses one score category for another, capturing its scoring tendencies and potential biases:
@@
2.2.2 Objective and Optimization
Objective.
Based on the model construction above, the optimization objective is to maximize the log conditional likelihood of the observed scoring labels $\mathbf{Y}=\{y^{(i,j;j^{\prime})}\mid 1\leq i\leq I,\ 1\leq j\leq J,\ 1\leq j^{\prime}\leq J\}$ contributed by $J$ LLM judges, i.e., $\log p(\mathbf{Y};\Theta)$ w.r.t. the parameters $\Theta=\{\bm{\alpha},\bm{\Pi}^{(1)},\ldots,\bm{\Pi}^{(J)}\}$.
Optimization.
In brief, as with most latent variable models (Bishop & Nasrabadi, 2006; Dawid & Skene, 1979), we apply the Expectation-Maximization (EM) algorithm (Dempter, 1977) to solve the optimization problem. [^2] We provide the detailed derivations in Appendix. First, the log-likelihood can be written as:
@@ 2.2.2 Objective and Optimization
where $\mathbf{y}^{(i,j)}=\{y^{(i,j;j^{\prime})}\mid 1\leq j^{\prime}\leq J\}$ denotes the set of scores assigned to response $\mathbf{r}^{(i,j)}$ by the $J$ models. Then, we use Jensen’s inequality (Bishop & Nasrabadi, 2006) to derive the Evidence Lower Bound (ELBO):
@@
where $q(t^{(i,j)})$ is a discrete distribution over the variable $t^{(i,j)}$. In the following, we then proceed to apply the general EM recipe to perform iterative calculations (concerning E-step and M-step) to solve the optimization problem $\Theta:=\underset{\Theta}{\operatorname{argmax}}\log p(\mathbf{Y};\Theta)$.
E-step (inference).
@@
M-step (learning).
Furthermore, by maximizing optimization objective in Equation 4 and using the standard Lagrange multiplier method (Bishop & Nasrabadi, 2006), we can obtain the closed-form solution for $\bm{\alpha}=\{\alpha^{(k)}\mid 1\leq k\leq K\}$ in Equation 7 shown below; and by equating the gradient of Equation 4 to zero, we can obtain the closed-form solution for $\{\bm{\Pi}^{(j^{\prime})}\}_{j^{\prime}=1}^{J}$ in Equation 8 shown below.
@@
$S(t^{(i,j)})=\sum_{k=1}^{K}q(t^{(i,j)}=k)\cdot s_{k},$ (9)
where $s_{k}$ denotes the score value corresponding to the $k$-th scoring category. [^3] For example, suppose that for a given response, we have $q(t^{(i,j)}=4)=0.5$, $q(t^{(i,j)}=5)=0.5$, along with the score values $s_{4}=4.0$, $s_{5}=5.0$. Then, the final aggregated score is $S(t^{(i,j)})=0.5\cdot 4.0+0.5\cdot 5.0=4.5$.
@@ 2.3 Selecting the Best
$\mathbf{r}_{\text{ensemble}}^{(i)}=\underset{\mathbf{r}^{(i,j)}}{\operatorname{argmax}}\{S(t^{(i,j)})\mid 1\leq j\leq J\},$ (10)
@@
The overall procedure for our proposed LLM-PeerReview is summarized in Algorithm 1.
Algorithm 1 LLM-PeerReview-Weighted
Input: Data $\mathcal{D}=\left\{\mathbf{x}^{(i)},\mathbf{R}^{(i)}\right\}_{i=1}^{I}$, where for each query $\mathbf{x}^{(i)}$, we have responses $\mathbf{R}^{(i)}=\{\mathbf{r}^{(i,j)}\mid 1\leq j\leq J\}$ from heterogeneous LLMs $\left\{\mathcal{M}_{j}\right\}_{j=1}^{J}$

### v2

@@ 2.2.2 Reasoning for LLM-PeerReview-W
where the distribution is parametrized by $\bm{\alpha}$. Next, similar to the concept of confusion matrix commonly used in machine learning (Bishop and Nasrabadi, 2006), we introduce an annotator-specific transition matrix $\bm{\Pi}^{(j^{\prime})}$ to model the probability that an LLM confuses one score category for another, capturing its scoring tendencies and potential biases:
@@
Figure 2: Probabilistic graphical representation.
Objective and optimization.
Based on the model construction above, the optimization objective is to maximize the log conditional likelihood of the observed scoring labels $\mathbf{Y}=\{y^{(i,j;j^{\prime})}\mid 1\leq i\leq I,\ 1\leq j\leq J,\ 1\leq j^{\prime}\leq J\}$ contributed by $J$ LLM judges, i.e., $\log p(\mathbf{Y};\Theta)$ w.r.t. the parameters $\Theta=\{\bm{\alpha},\bm{\Pi}^{(1)},\ldots,\bm{\Pi}^{(J)}\}$. As with most latent variable models, we apply the Expectation-Maximization (EM) algorithm (Dempter, 1977) to solve the optimization problem. [^3] We provide the detailed derivations in Appendix. First, the log-likelihood can be written as:
@@
where $\mathbf{y}^{(i,j)}=\{y^{(i,j;j^{\prime})}\mid 1\leq j^{\prime}\leq J\}$ denotes the set of scores assigned to response $\mathbf{r}^{(i,j)}$ by the $J$ models. Then, we use Jensen’s inequality (Bishop and Nasrabadi, 2006) to derive the Evidence Lower Bound (ELBO):
@@
where $q(t^{(i,j)})$ is a discrete distribution over the variable $t^{(i,j)}$. In the following, we then proceed to apply the general EM recipe to perform iterative calculations.
Optimization: E-step (inference).
@@
Optimization: M-step (learning).
Furthermore, by maximizing optimization objective in Equation 5 and using the standard Lagrange multiplier method (Bishop and Nasrabadi, 2006), we can obtain the closed-form solution for $\bm{\alpha}=\{\alpha^{(k)}\mid 1\leq k\leq K\}$ in Equation 8 shown below; and by equating the gradient of Equation 5 to zero, we can obtain the closed-form solution for $\{\bm{\Pi}^{(j^{\prime})}\}_{j^{\prime}=1}^{J}$ in Equation 9.
@@
$Score(\mathbf{r}^{(i,j)})=\sum_{k=1}^{K}q(t^{(i,j)}=k)\cdot s_{k},$ (10)
where $s_{k}$ denotes the score value corresponding to the $k$-th scoring category. [^4] For example, suppose that for a given response, we have $q(t^{(i,j)}=4)=0.5$, $q(t^{(i,j)}=5)=0.5$, along with the score values $s_{4}=4.0$, $s_{5}=5.0$. Then, the final aggregated score is $Score^{(i,j)}=0.5\cdot 4.0+0.5\cdot 5.0=4.5$.
@@ 2.3 Selecting the Best
$\mathbf{r}_{\text{ensemble}}^{(i)}=\underset{\mathbf{r}^{(i,j)}}{\operatorname{argmax}}\{Score(\mathbf{r}^{(i,j)})\mid 1\leq j\leq J\},$ (11)
@@
In summary, the overall procedure for our proposed LLM-PeerReview is provided in Algorithm 1.
Algorithm 1 LLM-PeerReview
Input: Data $\mathcal{D}=\{\mathbf{x}^{(i)},\mathbf{R}^{(i)}\}_{i=1}^{I}$, where for each query $\mathbf{x}^{(i)}$, we have responses $\mathbf{R}^{(i)}=\{\mathbf{r}^{(i,j)}\mid 1\leq j\leq J\}$ from heterogeneous LLMs $\left\{\mathcal{M}_{j}\right\}_{j=1}^{J}$